OpenStudy (anonymous):

find all the solutions to cos theta =1/2.

OpenStudy (anonymous):

I think \(\large \mathbb {cos~\theta =\frac{1}{2}}\rightarrow \theta =cos^{-1}(\frac{1}{2})=\pm \frac{\pi}{3}+k2\pi\)

OpenStudy (anonymous):

cos can only positive in the 1st and 4th so take the inverse for cos to get the angle this would for the 1st

OpenStudy (anonymous):

are there 2 solutions?

OpenStudy (anonymous):

Are the solutions in any interval?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

[0,360)

OpenStudy (anonymous):

Yes, as AJ01 said, the angles will be in 1st and 4th quadrant.

OpenStudy (anonymous):

first angle =60 second = 360-60=300

OpenStudy (anonymous):

good thats what i got! {60,300} is how I would write the answer correct?

OpenStudy (anonymous):

Yes, that's how we write the solution set.

OpenStudy (anonymous):

thanks!